OpenStudy (anonymous):

how do you solve reciprocal trigonometric functions for example:sec of 0 degrees, 30degrees, 45 degrees, 60 degrees , 90 degrees , 180 degrees, 270 degrees, 360 degrees

OpenStudy (anonymous):

hey, how's it going?

OpenStudy (bahrom7893):

just rewrite them as: Sec 0 = 1/Cos0 = 1 Csc 0 = 1/Sin0 = infinity

OpenStudy (anonymous):

what about the cot and csc of those angles as well

OpenStudy (bahrom7893):

Tan = Sin/Cos Cot = Cos/Sin Sec = 1/Cos Csc = 1/Sin

OpenStudy (bahrom7893):

now just plug the Sins and Coss in there and thats it..